Find the value of y in the formula y=5a-3b+c, if a=12, b=4, and c=8?
Answer by PRMath(122)

Find the value of y in the formula y=5a-3b+c, if a=12, b=4, and c=8?

If I understand your problem correctly, then all you have to do is "plug" the info you are given into your equation.

So, you have to remember that a = 12 b = 4 and c = 8

Now let's plug in:

y=5a-3b+c
............a...........b......c (see where we have plugged in the values for a, b and c?)
y = 5(12) - 3(4) + 8
y = 60 - 12 + 8
y= 56